OpenStudy (mpj4):

Calculus: can you use tangent half angle substitution on something like sin^2(x)?

OpenStudy (mpj4):

or it can only be used on non squared trig functs?

Parth (parthkohli):

You definitely can, but it is not worth it. The proper way to integrate is to use the identity:\[\sin^2 (x) = \dfrac{1 - \cos(2x)}{2}\]

OpenStudy (mpj4):

Ah, that helps a lot. Thank you!

Parth (parthkohli):

No problem. :)